OpenStudy (anonymous):

If 8x + 5x + 2x + 4x = 114, the 5x + 3 = (A) 12 (B) 25 (C) 33 (D) 47

OpenStudy (mayankdevnani):

8x + 5x + 2x + 4x = 114 19x=144 x=114/19=6

OpenStudy (theviper):

8x + 5x + 2x + 4x =114 19x=114 x=114/19=6 Now put the value in the equation 5x+3 5(6)+3 =30+3=33 Therefore C is the right answer :)

OpenStudy (anonymous):

sure?

OpenStudy (mayankdevnani):

so 5*6+3=?

OpenStudy (anonymous):

33

OpenStudy (mayankdevnani):

can you solve it?

OpenStudy (theviper):

Fully sure @rambo2210 :D

OpenStudy (mayankdevnani):

ya @rambo2210

OpenStudy (anonymous):

33 is the answer right?

OpenStudy (mayankdevnani):

yup

OpenStudy (theviper):

Don't be confuse Obviously 33 is the right answer :)